OpenStudy (anonymous):

x=8√x solve for real solutions

OpenStudy (anonymous):

x=0,64 are the solutions

OpenStudy (anonymous):

I have the solutions, how do you get them?

OpenStudy (anonymous):

square both sides x^2 = 64x x(x-64) = 0 x =0 or 64

OpenStudy (anonymous):

^ That's how you get them.

OpenStudy (anonymous):

^4√x^2+16=√5 solve for real solutions? answer is 4 but don't get how to get that